OpenStudy (anonymous):

if f(x)=12x pwr 5 and g(x)= -3x pwr 2, determine (f.g)(-1) [given: (f.g) (x) =f(x).g(x)] plz help..!

OpenStudy (luffingsails):

dyanesh, If I read what you wrote correctly, you are given: f*g(x) = f(x)*g(x) And you are asked to solve f*g(-1). So, it is really a plug and play problem. f*g(-1) = f(-1)*g(-1) So, by substitution, \[f*g(-1)=12(-1)^{5} * (-3)(-1)^{2}\] \[f*g(-1) = 12(-1) * (-3)(1) = -12 * -3 = 36\] Hope that helps.